OpenStudy (anonymous):

can someone help me on 4 multiple questions on algebraic postulates?

OpenStudy (anonymous):

Select the postulate that is illustrated for the real numbers. 3x + 3 = 3(x + 1) The commutative postulate for multiplication Multiplication by one The addition inverse postulate Commutative postulate for addition The distributive postulate The addition of zero postulate The multiplication inverse

OpenStudy (anonymous):

Select the postulate of equality or inequality that is illustrated. 3 + 2 5 and 3 + 2 = 5 are not both true. addition postulate of inequality the multiplication postulate of inequality the symmetric postulate of equality the transitive postulate of equality comparison postulate the reflexive postulate of equality the transitive postulate of inequality

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2013/1/49.gif Select the postulate of equality or inequality that is illustrated. addition postulate of inequality the multiplication postulate of inequality the symmetric postulate of equality the transitive postulate of equality comparison postulate the reflexive postulate of equality the transitive postulate of inequality

OpenStudy (anonymous):

nevermind i found it all out myself

OpenStudy (anonymous):

what was the answer to Select the postulate of equality or inequality that is illustrated. addition postulate of inequality the multiplication postulate of inequality the symmetric postulate of equality the transitive postulate of equality comparison postulate the reflexive postulate of equality the transitive postulate of inequality @Morganaemmalynn